OpenStudy (anonymous):

evaluate 183 - x to 2nd power when x = 5 183-x 2nd power simplify and type and integer or a decimal is the ans. 31684

OpenStudy (anonymous):

evaluate 183 - x to 2nd power when x = 5 183-x 2nd power simplify and type and integer or a decimal

OpenStudy (anonymous):

(183-x)^2=178^2= 31684 =3.1684*10^4